OpenStudy (anonymous):

Let f(x)=x^sinx For x>0 Find f'(x)

OpenStudy (kira_yamato):

Let y = f(x) y = x^sinx ln y = sin x ln x y'/y = (sin x)/x + cos x ln x y' = x^sin x [(sin x)/x + cos x ln x] f'(x) = y' = x^sin x [(sin x)/x + cos x ln x]

OpenStudy (anonymous):

Im still confused, i dont know what properties to use.

OpenStudy (agent0smith):

Log properties. You can also change the method in the third step ln y = sin x ln x Put both sides to the power of e y = e^(sin x ln x) then differentiate from here if you prefer.

OpenStudy (anonymous):

So i see why you ln them. But how did the y turn into y'?

OpenStudy (anonymous):

Im having lots of trouble in indirect differentiation.

OpenStudy (anonymous):

Implicit

OpenStudy (agent0smith):

Derivative of \[\large \frac{ d }{ dx } \ln f(x) = \frac{ f'(x) }{ f(x) }\]It's no different with y. \[\large \frac{ d }{ dx } \ln y = \frac{ y' }{ y}\]since the derivative of y (w.r.t. x) is y'

OpenStudy (anonymous):

Alright thanks